OpenStudy (anguyennn):

how do solve 3x^3+2x^2-12x-8=0

zepdrix (zepdrix):

Ooo I think this one is going to work out nicely with grouping! :)

OpenStudy (anguyennn):

how do you do that?

OpenStudy (anguyennn):

yes that is what the question is asking me to do actually haha :)

zepdrix (zepdrix):

Oh interesting c: hah

zepdrix (zepdrix):

\[\large\rm \color{#DD4747}{3x^3-12x}+\color{#3366CF}{2x^2-8}=0\]Notice that the cubic term and the first degree term are both divisible by 3, while the squared term and constant term are both even, divisible by 2, so let's group them like this!

zepdrix (zepdrix):

We want to do some factoring. What can we factor out of the red guys? What do they have in common?

OpenStudy (anguyennn):

\[\large\rm \color{#DD4747}{3x^3-12x}+\color{#3366CF}{2x^2-8}=0\]

OpenStudy (anguyennn):

all i see is that

OpenStudy (anguyennn):

I don't see any equations

zepdrix (zepdrix):

The equation didn't show up for you? Are you using Internet Explorer or some junk like that? :c lol

OpenStudy (anguyennn):

I am using safari

zepdrix (zepdrix):

oh boy :( hmm

zepdrix (zepdrix):

Reload browser? :o Maybe the plugin just crashed for you or something

OpenStudy (anguyennn):

ok please hold on ok

OpenStudy (anguyennn):

oh yes i see it now!!

zepdrix (zepdrix):

yay c:

OpenStudy (anguyennn):

so it would be (3x+2) (x^2-4)

zepdrix (zepdrix):

Oh, you did it all the way already XD haha nice!

OpenStudy (anguyennn):

so would the final answer be x=2/3, -2, 2?

zepdrix (zepdrix):

-2/3, ya? :) Gotta subtract the 2 over

OpenStudy (anguyennn):

oh!!!

OpenStudy (anguyennn):

but wait, what about x^3-8x^2=0

zepdrix (zepdrix):

Oh this is another problem? I see :)

OpenStudy (anguyennn):

this is another one

OpenStudy (anguyennn):

yes

zepdrix (zepdrix):

So they each have at least two x's, multiplying, ya? So we can pull an x^2 out of each.

zepdrix (zepdrix):

\[\large\rm x^2(x-8)=0\]Something like that, ya? :d

OpenStudy (anguyennn):

yes

zepdrix (zepdrix):

Apply your `Zero-Factor Property` again :) \(\large\rm x^2=0\) and \(\large\rm x-8=0\) and solve for x in each case.

OpenStudy (anguyennn):

I know one of them would be x=8

OpenStudy (anguyennn):

but what about for the x^2? that is what i am confused about

zepdrix (zepdrix):

\(\large\rm x^2=0\) Take square root, \(\large\rm \pm x=0\) Leading to, \(\large\rm x=0\)

zepdrix (zepdrix):

square root of zero is... zero :) ya?

OpenStudy (anguyennn):

yes

OpenStudy (anguyennn):

but when you square root x it becomes 10

zepdrix (zepdrix):

?

OpenStudy (anguyennn):

oh wait never mind, i am talking nonsense

zepdrix (zepdrix):

x is being squared. we perform the `opposite` operation of squaring to remove the square. square root is the opposite of squaring. So when we apply the square root, they `undo` one another.

OpenStudy (anguyennn):

wow thanks I understand it now!!

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anguyennn):

are you able to help me with some other questions since you are here already :P

OpenStudy (anguyennn):

you are a really great help btw thank you i appreciate it

zepdrix (zepdrix):

uh sure +_+

zepdrix (zepdrix):

you can post it, i gotta make some food real quick :D brb

OpenStudy (anguyennn):

I am not sure how to do c, f, or g.

zepdrix (zepdrix):

What did you get for b? Your profit function.

OpenStudy (anguyennn):

I got 13.83n-102.4

OpenStudy (anguyennn):

I showed you a picture as well

OpenStudy (anguyennn):

the second document

zepdrix (zepdrix):

oh lol :)

OpenStudy (anguyennn):

yup :)

zepdrix (zepdrix):

Since this is a `linear function`, \(\large\rm \text{rate of change = slope}\) So the slope of your P(n) function is your rate of change.

OpenStudy (anguyennn):

how do you find out the slope? mx+b?

zepdrix (zepdrix):

Yes, m :)

zepdrix (zepdrix):

We can write the number like this if it makes things easier:\[\large\rm m=\frac{13.83}{1}\]Remember that slope is `rise over run`. So I would like two separate numerical values in my slope so I can interpret it. The 13.83 represents profit, the 1 represents 1 pizza sold. So this value represents the amount of profit made per pizza!

OpenStudy (anguyennn):

hmm

OpenStudy (anguyennn):

so would 1 be the y coordinate on the graph?

zepdrix (zepdrix):

slope is `change in y` divided by `change in x`. So the bottom number is the x ( n in this case ).

OpenStudy (anguyennn):

oh ok thank you so much for your help!

OpenStudy (anguyennn):

until next time :)

zepdrix (zepdrix):

`practical domain` You don't want to open a pizza shop and sell two pizzas every day. You'll end up losing money because of your operating costs! So you want your domain to start from the `break-even point`.

OpenStudy (anguyennn):

which is going to be n=7.40?